Factor completely.
-3x^3 - 6x^2 + 189x

Answer by mananth(12270)

-3x^3 - 6x^2 + 189x
-3x(x^2+2x-63)=0
-3x(x^2+9x-7x-63)=0
-3x(x(x+9)-7(x+9))=0
-3x(x-7)(x+9)=0
(x-7)(x+9)=0/-3x
(x-7)(x+9)=0
x=7 or-9